OpenStudy (anonymous):

What are the local extremes of y=x^2-x-1

OpenStudy (anonymous):

you need to find the vertex i think

OpenStudy (anonymous):

if you know calculus, differentiate the RHS, 2x-1=0, so x=1/2. extremes = (0.5,-1.25)

OpenStudy (anonymous):

You don't have to. You can find the vertex by complete the square